# Improving RL Exploration for LLM Reasoning through Retrospective Replay

### v1

@@ 3.3 Discussion
Sensitivity analysis. Compared to the vanilla PPO, we introduce an additional hyper-parameter, the replay coefficient $p$. During the sampling process, the policy has a probability of $p$ to generate a completion from a state in the buffer and a probability of $1-p$ to directly generate a solution from the problem. To effectively investigate the impact of the replay coefficient, we set it to six different values (i.e., 0, 0.1, 0.2, 0.3, 0.4, and 0.5) and conducted experiments. Figure 3 shows that, when the selection probability is set to 0.4, the model achieves optimal performance. Moreover, it shows that RRL under different replay coefficients can also bring improvements compared to the vanilla PPO, indicating our method can both achieve better performance under a range of settings.
@@ 4 Related Work
Exploration is a critical concept in RL [27, 37, 6, 23, 21, 1]. When faced with complex reasoning problems, large language models need to explore a vast generation space to identify correct solution paths and learn from these experiences [9, 45, 20, 28, 17]. Pre-trained language models often cannot be directly used for reinforcement learning because their exploration space is too large, which negatively impacts exploration efficiency [13, 26]. Instruction fine-tuning can be considered a method to enhance exploration efficiency in reinforcement learning, as it narrows the exploration space of LLMs to a subspace, significantly reducing the complexity of exploration [40, 46, 10]. Moreover, some studies have demonstrated that few-shot-based in-context learning can also enable LLMs to follow instructions, thereby improving exploration efficiency in RL [13, 41].
@@
A similar but entirely different concept is experience replay, which involves storing the experiences including states and actions in a buffer and replaying them to update policy and value functions [30, 44]. In contrast, RRL replays previously explored potential intermediate states to continue exploration. In addition, some approaches also introduce curriculum reinforcement learning, allowing models to start learning from simpler intermediate nodes and progressively tackle complete complex problems [42, 9, 12, 36]. Compared to these methods, our approach aims to maintain a high level of exploration throughout the training process. We are the first to reveal that LLMs, during exploration, may regrettably forget previously explored but not fully pursued potential solution ideas (i.e., states) due to policy gradients.

### v2

@@ 3.3 Discussion
Sensitivity analysis. Compared to the vanilla PPO, we introduce an additional hyper-parameter, the replay coefficient $p$. During the sampling process, the policy has a probability of $p$ to generate a completion from a state in the buffer and a probability of $1-p$ to directly generate a solution from the problem. To effectively investigate the impact of the replay coefficient, we set it to six different values (i.e., 0, 0.1, 0.2, 0.3, 0.4, and 0.5) and conducted experiments. Figure 3 shows that, when the selection probability is set to 0.1, the model achieves optimal performance. Moreover, it shows that RRL under different replay coefficients can also bring improvements compared to the vanilla PPO, indicating our method can both achieve better performance under a range of settings.
@@ 4 Related Work
Recently, large language models have shown remarkable ability in reasoning and general conversation tasks [6, 47, 27, 4, 19, 15, 14, 5, 13, 10, 53]. Reinforcement learning has gradually become a core technology for enhancing the capabilities of large language models [49, 40, 38, 17]. Exploration is a critical concept in reinforcement learning [32, 44, 8, 26, 1, 11, 50]. When faced with complex reasoning problems, large language models need to explore a vast generation space to identify correct solution paths and learn from these experiences [11, 55, 25, 33, 22]. Pre-trained language models often cannot be directly used for reinforcement learning because their exploration space is too large, which negatively impacts exploration efficiency [18, 31]. Instruction fine-tuning can be considered a method to enhance exploration efficiency in reinforcement learning, as it narrows the exploration space of LLMs to a subspace, significantly reducing the complexity of exploration [47, 56, 12]. Moreover, some studies have demonstrated that few-shot-based in-context learning can also enable LLMs to follow instructions, thereby improving exploration efficiency in RL [18, 48].
@@
A similar but entirely different concept is experience replay, which involves storing the experiences including states and actions in a buffer and replaying them to update policy and value functions [35, 54]. In contrast, RRL replays previously explored potential intermediate states to continue exploration. In addition, some approaches also introduce curriculum reinforcement learning, allowing models to start learning from simpler intermediate nodes and progressively tackle complete complex problems [51, 11, 43]. Compared to these methods, our approach aims to maintain a high level of exploration throughout the training process. We are the first to reveal that LLMs, during exploration, may regrettably forget previously explored but not fully pursued potential solution ideas (i.e., states) due to policy gradients.
@@ 6 Conclusion
Acknowledgments
The authors wish to thank the anonymous reviewers for their helpful comments. This work was partially funded by Guangdong S&T Program 2024B0101050003.